OpenStudy (anonymous):

gloria had rectangle garden plot last year with an area of 60 feet square.This year glorias garden plot is one foot wider and three feet shorter than last years garden,but it has the same area.What were the dimensions of the garden last year?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mary.rojas

OpenStudy (anonymous):

@mathmath333

OpenStudy (mary.rojas):

I think it is something like...... 60= (1+w)*(L-3) Where w is the width and L is the length You can use any letters to represent them in the equation. Now solve for w and then solve for L. And that will be the answer...I think Good luck and hope it helps

OpenStudy (anonymous):

thank you my teacher gave me this equation he said the A=l times w so he said 30*2

OpenStudy (mary.rojas):

hmmmm...but you may have to keep in mind that this is for the new garden...so you would have to keep in mind that teh original was w*L=60

OpenStudy (anonymous):

thanks for the help

OpenStudy (mary.rojas):

Well, that is good that your teacher helped you! :)

OpenStudy (anonymous):

h:)

OpenStudy (anonymous):

@Bush993

OpenStudy (anonymous):

@Bush993

OpenStudy (anonymous):

@Bush993

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

i could explain this to you but i'll take a while this website is awesome thou http://www.algebra.com/algebra/homework/word/geometry/Geometry_Word_Problems.faq.question.438542.html

OpenStudy (anonymous):

thank you @superhelp101

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

as mary.rojas has started already, solve : WL = 60 (W+1)(L-3)=60

OpenStudy (anonymous):

60=W*L W=60/L .....1 60=(W+1)(L-3) 60=(60/L+1)(L-3) 60=60-180/L+L-3 60=57-180/L+L 60L=57L-180+L^2 3L-L^2=180 L^2-3L-180=0 (L+12)(L-15)=0 L=-12 L=15 then L=15 W=60/15 then W=4 in the last year W+1=5 L-3=12

OpenStudy (anonymous):

@ganeshie8 i dont understand how to do this

OpenStudy (anonymous):

can you plz help me through the equation above @ganeshie8

ganeshie8 (ganeshie8):

WL = 60 (W+1)(L-3)=60

ganeshie8 (ganeshie8):

from the first equation, L = 60/W plug this in second equation, you get : (W+1)(60/W - 3) = 60

ganeshie8 (ganeshie8):

expand that, you will get a quadratic.. solve it

OpenStudy (anonymous):

um -5 and 4 is w i

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

what is a quadratic @ganeshie8

ganeshie8 (ganeshie8):

thats right! pick W = 4 because the lengths cannot be negative http://www.wolframalpha.com/input/?i=solve+W*L+%3D+60%2C+%28W%2B1%29%28L-3%29%3D60%2CW%3E0

OpenStudy (anonymous):

whats the last year and this year length though

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Isaiah.Feynman can you walk me through this problem

OpenStudy (isaiah.feynman):